# Resolvent4py: a parallel Python package for analysis, model reduction and control of large-scale linear systems

### v1

@@ Abstract
Python , Parallel computing , Petsc4py , Slepc4py , Mpi4py , Model reduction , Resolvent analysis , Stability analysis , Harmonic resolvent analysis
@@ 1 Motivation and significance
The development of this package is motivated by recent (and not-so-recent) advances in the field of fluid mechanics, where linear analysis has been shown to provide great insights into the flow physics. For example, linear systems theory can be used to study the stability of fluid flows around equilibria, develop feedback control strategies to modify the flow behavior and compute reduced-order models to accelerate physical simulations. For systems of moderate dimension (e.g., fewer than 10,000 states), all these tasks can be performed straightforwardly with a few lines of Python code thanks to scientific computing libraries like scipy. For larger systems, the need for MPI-based parallelism significantly increases the implementation complexity of these algorithms, and it often requires the development of application-specific software written in compiled languages such as C, C++ and Fortran. The objective of this package is to provide a user-friendly pythonic environment similar to that of scipy and numpy, while exploiting the large-scale parallelism offered by the mpi4py, petsc4py and slepc4py libraries. In turn, this enables the development of user-friendly, streamlined and efficient code to perform day-to-day engineering tasks and physical analyses of large-scale systems.
@@ 2 Software description
resolvent4py is a Python library for the analysis, model reduction and control of large-scale linear systems. It relies on mpi4py [1] for multi-processing parallelism, and it leverages the data structures and functionalities provided by the petsc4py [2] and slepc4py [3] libraries.
@@ 2.1 Software architecture
, respectively), the action of the inverse $L^{-1}$ on vectors, solve(), etcetera. Below is a list of concrete subclasses of LinearOperator that are provided by resolvent4py:
@@
    where $U$ and $V$ are tall and skinny matrices stored as MPI-distributed SLEPc BV structures [^1] The SLEPc BV structure (where BV stands for “Basis Vector”) holds a tall and skinny dense matrix whose rows are distributed across different MPI processors. The columns, however, are not distributed, and this allows for significantly reduced parallel overhead when the number of MPI processors is much larger than the number of columns, as is usually the case in practice. , while $\Sigma$ is a sequential two-dimensional numpy array of conformal size.
@@
    LowRankUpdatedLinearOperator: the linear operator $L$ is defined as $L=A+M$, where $A$ is itself a resolvent4py linear operator and $M$ is a low-rank linear operator, as described in the previous bullet point. An operator of this form arises in several applications and, if $A$ and $L$ are full-rank, then the inverse $L^{-1}$ can be computed using the Woodbury matrix inversion lemma [4],
@@ 2.2 Software functionalities
    Eigendecomposition: the user can compute the left and right eigedecomposition of the linear operator $L$ using the Arnoldi algorithm [5]. The shift-and-invert technique can be used to compute the eigenvalues that are closest to a target value $s\in\mathbb{C}$, where $\mathbb{C}$ is the set of complex numbers.
@@
    Linear time-invariant balanced truncation: given the linear operator $L$, and appropriately-defined input and output matrices $B$ and $C$, the user may compute and balance the associated controllability and observability Gramians for model reduction purposes [17, 18, 19]. Specifically, we implement the algorithm outlined in [20].
@@ 2.3 Sample code snippet
L = res4py.linear_operators.MatrixLinearOperator(A, ksp)
@@ 3.1 Hypersonic flow over a compression ramp
The first example is a Mach–$7.7$ flow over a $15\deg$ compression ramp of length $L=0.1\,m$. In the freestream, the Mach number is $M_{\infty}=7.7$, the pressure is $p_{\infty}=760\,Pa$, the temperature is $T_{\infty}=125\,K$, the streamwise velocity is $u_{\infty}=1726\,m/s$ and the Reynolds number based on $L$ is $Re_{\infty,L}=4.2\times 10^{5}$. This flow configuration is studied extensively in [21], and we refer to the latter for details on boundary conditions, spatial discretization, mesh and non-dimensionalization. As discussed in [21], this flow admits a two-dimensional steady state solution (see figure 1) that is unstable with respect to three-dimensional perturbations. These instabilities can be detected by linearizing the governing equations about the aforementioned steady state and studying the dynamics of three-dimensional, spanwise-periodic perturbations with wavenumber $\beta$. Upon linearization, the spatially-discretized compressible Navier-Stokes equations can be written compactly as
@@
where the state vector $q_{\beta}$ denotes the wavenumber-$\beta$ Fourier coefficients of the spatially-discretized flow variables. Here, given a computational grid of size $n_{y}\times n_{x}=236\times 1076$ and $5$ flow variables, the size of the state vector is $N=5\,n_{y}\,n_{x}\approx 1.27\times 10^{6}$. The eigenvalues of $A_{\beta}$ and corresponding eigenvectors for (non-dimensional) wavenumber $\beta=2\pi/0.066$ are computed using resolvent4py with $240$ processors on Stempede3 at Texas Advanced Computing Center. The eigenvalues are shown in figure 2, where we see good agreement with the results in Cao et al. [21]. Indeed, we find that the flow exhibits multiple instabilities corresponding to the five eigenvalues $\lambda$ with real part greater than zero. The flow structures associated with these instability modes are shown in figure 3, where we plot contours of the real part of the spanwise velocity component from the eigenvectors corresponding to three unstable eigenvalues. Here, we find that the eigenvector corresponding to $\lambda_{1}$ exhibits very good qualitative agreement with its counterpart in figure 14a in [21]. The eigenvectors corresponding to $\lambda_{2}$ and $\lambda_{3}$ are very similar in nature to those in figures 14b and 14c in [21], but it is important to observe that a direct comparison cannot be made since the modes shown in figures 14b and 14c in [21] correspond to wavenumbers $\beta=0.070$ and $\beta=0.079$, respectively.
@@ 3.2 Hypersonic flow over a cone
Another illustrative example is a Mach-5.9 flow over a 7 deg half-angle blunt cone. The freestream quantities are $M_{\infty}=5.9$, $p_{\infty}=3396.3$ $Pa$, $T_{\infty}=76.74$ $K$, $u_{\infty}=1036.1$ $m/s$, and $Re_{\infty}=18\times 10^{6}$ and the domain length is $L=0.4893\,m$. Bluntness effects produce unique instability phenomena in comparison with sharp cones, and are extensively studied using various DNS and operator theoretic methods in receptivity studies  [22]. At high Mach numbers, a discrete higher-order mode termed the second Mack mode becomes destabilized and dominates laminar-to-turbulent transition. This mechanism manifests itself as acoustic waves trapped within the boundary layer. Likewise, response modes are modulated by the entropy layer observed near the front of the cone, which can contribute to distinct turbulent transition mechanisms.
@@
The spatially-discretized Navier-Stokes equations are linearized about the base flow shown in figure 4, and the singular value decomposition of the resolvent operator is computed using the RSVD-$\Delta$t algorithm using a script almost identical to the one shown in section 2.3. The temperature perturbations extracted from the leading right and left singular vectors of the resolvent are seen in figure  5. The response mode corresponding to maximal amplification appears trapped in the boundary layer, while the forcing mode follows the decaying entropy layer’s contour within the boundary layer, indicative of the importance of the high entropy gradients of the flow due to bluntness.
Figure 5: Temperature forcing and response modes computed using RSVD-$\Delta$t plotted in streamwise and wall-normal coordinates normalized by the streamwise length of the domain.
@@ 4 Impact
Additionally, we believe that the current version of resolvent4py can serve as a stepping stone for additional software development—both by the original developers and the scientific community at large—that will be featured in future releases. For example, the infrastructure of resolvent4py enables the straightforward implementation of time-periodic balanced truncation using “Frequential Gramians” [28], harmonic resolvent analysis via time stepping [29], wavelet-based resolvent analysis [30] and One-Way Navier-Stokes (OWNS) for slowly-developing flows [31].
@@ 5 Conclusions
In this manuscript, we introduced resolvent4py, a parallel Python package for the analysis, model reduction and control of large-scale linear systems. Large-scale linear systems with millions or billions of degrees of freedom are ubiquitous in the sciences and engineering, and this package is built to provide a friendly pythonic environment for rapid development and deployment of MPI-based parallel software to solve problems of practical interest. Our library leverages mpi4py for MPI-based parallelism, and it uses the functionalities and data structures provided by petsc4py and slepc4py to enable stability analysis, input/output analysis, model reduction, control and optimization of linear time-invariant and time-periodic systems. Additionally, the package ships with several functions and features that are available to the user through the resolvent4py namespace and that facilitate user-specific and application-specific software development within the larger resolvent4py infrastructure.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2139536, issued to the University of Illinois at Urbana-Champaign by the Texas Advanced Computing Center under subaward UTAUS-SUB00000545 with Dr. Daniel Stanzione as the PI. DB gratefully acknowledges support from the Office of Naval Research (N00014-21-1-2256), and CR gratefully acknowledges support from the Air Force Office of Scientific Research (FA9550-19-1-0005). TC and VA gratefully acknowledge support from The Boeing Company (CT-BA-GTA-1) and the Office of Naval Research (N00014-25-1-2072). BV was supported by the LDRD Program at Sandia National Laboratories. Sandia is managed and operated by NTESS under DOE NNSA contract DE-NA0003525. The computations in section 3.1 were performed on TACC’s Stampede3 under under ACCESS allocation CTS090004.

### v2

@@ Abstract
Python , Parallel computing , Model reduction , Resolvent analysis , Stability analysis , Harmonic resolvent analysis
@@ 1 Motivation and significance
The development of this package is motivated by recent (and not-so-recent) advances in the field of fluid mechanics, where linear analysis has been shown to provide great insights into flow physics. For example, linear systems theory can be used to study the stability of fluid flows around equilibria, develop feedback control strategies to modify the flow behavior and compute reduced-order models to accelerate physical simulations. (We refer the reader to the foundational work of [1] and to the review papers in [2, 3] for a more in-depth overview of the importance and impact of linear systems theory in the field of fluid mechanics.) For systems of moderate dimension (e.g., fewer than 10,000 states), all these tasks can be performed straightforwardly with a few lines of Python code thanks to scientific computing libraries like scipy. For larger systems, the need for distributed-memory parallelism significantly increases the implementation complexity of these algorithms, and it often requires the development of application-specific software written in compiled languages such as C, C++ and Fortran. In fact, until the recent development of Python libraries like mpi4py and petsc4py, compiled languages represented the only way to write MPI-based [^1] MPI stands for Message Passing Interface, and it refers to the communication protocol used to exchange data across different processors in distributed-memory systems. code. The objective of this package is to address these issues and provide a user-friendly pythonic environment similar to that of scipy and numpy, while exploiting the large-scale parallelism offered by the mpi4py, petsc4py and slepc4py libraries. In turn, this enables the development of user-friendly, streamlined and efficient code to perform day-to-day engineering tasks and physical analyses of large-scale systems.
@@ 2 Software description
resolvent4py is a Python library for the analysis, model reduction and control of large-scale linear systems. It relies on mpi4py [4] for distributed-memory parallelism, and it leverages the data structures and functionalities provided by the petsc4py [5] and slepc4py [6] libraries.
@@ 2.1 Software architecture
, respectively), the action of the inverse $L^{-1}$ on vectors, solve(), etc. Below is a list of concrete subclasses of LinearOperator that are provided by resolvent4py:
@@
    where $U$ and $V$ are “tall and skinny” matrices (i.e., matrices with much fewer columns than rows) stored as MPI-distributed SLEPc BV structures [^2] The SLEPc BV structure (where BV stands for “Basis Vector”) holds a tall and skinny dense matrix whose rows are distributed across different MPI processors. The columns, however, are not distributed, and this allows for significantly reduced parallel overhead when the number of MPI processors is much larger than the number of columns, as is usually the case in practice. , while $\Sigma$ is a sequential (i.e., non-distributed) two-dimensional numpy array of appropriate size.
@@
    LowRankUpdatedLinearOperator: the linear operator $L$ is defined as $L=A+M$, where $A$ is itself a resolvent4py linear operator and $M$ is a low-rank linear operator, as described in the previous bullet point. An operator of this form arises in several applications (e.g., in linear feedback control theory [7]) and, if $A$ and $L$ are full-rank, then the inverse $L^{-1}$ can be computed using the Woodbury matrix inversion lemma [8],
@@ 2.2 Software functionalities
    Eigendecomposition: the user can compute the left and right eigendecomposition of the linear operator $L$ using the Arnoldi algorithm [9]. The shift-and-invert technique [10] can be used to compute the eigenvalues that are closest to a target value $s\in\mathbb{C}$, where $\mathbb{C}$ is the set of complex numbers.
@@
    Linear time-invariant balanced truncation: given the linear operator $L$, and appropriately-defined input and output matrices $B$ and $C$, the user may compute and balance the associated controllability and observability Gramians for model reduction purposes [21, 22, 23, 24]. Specifically, we implement the algorithm outlined in [25]. We remark that there exist additional model reduction formulations for linear system, (see, e.g., [26, 27, 28, 29] for $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ model reduction and variations thereof), but these are not yet implemented in the package.
@@ 2.3 Sample code snippet
L = res4py.linear_operators.MatrixLinearOperator(A)
@@
Remark 2
We would like to stress how this snippet exposes the application-agnostic nature of our package. The governing equations (e.g., the linearized fluid dynamics equations, the linearized molecular dynamics equations, or any other linear equation from other branches of physics) are embedded in the sparse matrix $A$ that is passed as an input to the script. Once $A$ is loaded from file, resolvent4py performs the desired analysis without additional knowledge of the underlying physics.
@@ 3.1 Hypersonic flow over a compression ramp
The first example is a Mach–$7.7$ flow over a $15\deg$ compression ramp of length L=0.1mL=0.1\,m. In the freestream, the Mach number is $M_{\infty}=7.7$, the pressure is $p_{\infty}=760\,Pa$, the temperature is $T_{\infty}=125\,K$, the streamwise velocity is $u_{\infty}=1726\,m/s$ and the Reynolds number based on $L$ is $Re_{\infty,L}=4.2\times 10^{5}$. This flow configuration is studied extensively in [30], and we refer to the latter for details on boundary conditions, spatial discretization, mesh and non-dimensionalization. As discussed in [30], this flow admits a two-dimensional steady state solution (see figure 1) that is unstable with respect to three-dimensional perturbations. These instabilities can be detected by linearizing the governing equations about the aforementioned steady state and studying the dynamics of three-dimensional, spanwise-periodic perturbations with wavenumber $\beta$. Upon linearization, the spatially-discretized compressible Navier-Stokes equations can be written compactly as
@@
where the state vector $q_{\beta}$ denotes the wavenumber-$\beta$ Fourier coefficients of the spatially-discretized flow variables. Here, given a computational grid of size $n_{y}\times n_{x}=236\times 1076$ and $5$ flow variables, the size of the state vector is $N=5\,n_{y}\,n_{x}\approx 1.27\times 10^{6}$. The eigenvalues of $A_{\beta}$ and corresponding eigenvectors for (non-dimensional) wavenumber $\beta=2\pi/0.066$ are computed using resolvent4py with $240$ processors on Stampede3 at Texas Advanced Computing Center. The eigenvalues are shown in figure 2, where we see good agreement with the results in Cao et al. [30]. Indeed, we find that the flow exhibits multiple instabilities corresponding to the five eigenvalues $\lambda$ with real part greater than zero. The flow structures associated with these instability modes are shown in figure 3, where we plot contours of the real part of the spanwise velocity component from the eigenvectors corresponding to three unstable eigenvalues. Here, we find that the eigenvector corresponding to $\lambda_{1}$ exhibits very good qualitative agreement with its counterpart in figure 14a in [30]. The eigenvectors corresponding to $\lambda_{2}$ and $\lambda_{3}$ are very similar in nature to those in figures 14b and 14c in [30], but it is important to observe that a direct comparison cannot be made since the modes shown in figures 14b and 14c in [30] correspond to wavenumbers $\beta=0.070$ and $\beta=0.079$, respectively.
@@ 3.2 Hypersonic flow over a cone
Another illustrative example is a Mach-5.9 flow over a 7 deg half-angle blunt cone. The freestream quantities are $M_{\infty}=5.9$, $p_{\infty}=3396.3$ PaPa, $T_{\infty}=76.74$ $K$, $u_{\infty}=1036.1$ $m/s$, and $Re_{\infty}=18\times 10^{6}$ and the domain length is L=0.4893mL=0.4893\,m. Bluntness effects produce unique instability phenomena in comparison with sharp cones, and are extensively studied using various DNS and operator theoretic methods in receptivity studies  [31]. At high Mach numbers, a discrete higher-order mode termed the second Mack mode becomes destabilized and dominates laminar-to-turbulent transition. This mechanism manifests itself as acoustic waves trapped within the boundary layer. Likewise, response modes are modulated by the entropy layer observed near the front of the cone, which can contribute to distinct turbulent transition mechanisms.
@@
[…]
@@ 4 Impact
Additionally, we believe that the current version of resolvent4py can serve as a stepping stone for additional software development—both by the original developers and the scientific community at large—that will be featured in future releases. For example, the infrastructure of resolvent4py enables the straightforward implementation of time-periodic balanced truncation using “Frequential Gramians” [37], harmonic resolvent analysis via time stepping [38], wavelet-based resolvent analysis [39] and One-Way Navier-Stokes (OWNS) for slowly-developing flows [40]. Furthermore, it should be possible to implement the algorithms discussed in [41, 42] for the efficient computation of low-rank solutions to Lyapunov and Riccati equations of large-scale, sparse linear systems such as those that arise from the spatial discretization of partial differential equations. These solutions could then be leveraged for model reduction and closed-loop control design.
@@ 5 Conclusions
In this manuscript, we introduced resolvent4py, a parallel Python package for the analysis, model reduction and control of large-scale linear systems. Large-scale linear systems with millions or billions of degrees of freedom are ubiquitous in the sciences and engineering, and this package is built to provide a friendly pythonic environment for rapid development and deployment of MPI-based parallel software to solve problems of practical interest. Our library leverages mpi4py for distributed-memory parallelism, and it uses the functionalities and data structures provided by petsc4py and slepc4py to enable stability analysis, input/output analysis, model reduction, control and optimization of linear time-invariant and time-periodic systems. Additionally, the package ships with several functions and features that are available to the user through the resolvent4py namespace and that facilitate user-specific and application-specific software development within the larger resolvent4py infrastructure. Finally, although the primary focus of this package is on linear systems, we wish to remark that the data structures and functionalities of resolvent4py can be extended straightforwardly to accommodate the analysis, model reduction and control of large-scale nonlinear systems.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2139536, issued to the University of Illinois at Urbana-Champaign by the Texas Advanced Computing Center under subaward UTAUS-SUB00000545 with Dr. Daniel Stanzione as the PI. DB gratefully acknowledges support from the Office of Naval Research (N00014-21-1-2256), and CR gratefully acknowledges support from the Air Force Office of Scientific Research (FA9550-19-1-0005). TC and VA gratefully acknowledge support from The Boeing Company (CT-BA-GTA-1) and the Office of Naval Research (N00014-25-1-2072). BV was supported by the LDRD Program at Sandia National Laboratories. Sandia is managed and operated by NTESS under DOE NNSA contract DE-NA0003525. The computations in section 3.1 were performed on TACC’s Stampede3 under ACCESS allocation CTS090004.